# Adaptive Control for Triadic Human-Robot-FES Collaboration in Gait Rehabilitation: A Pilot Study

### 01 Feb 2024 v1

@@ I INTRODUCTION
To further enhance the outcomes of physiotherapy, the combination of robotic assistance with functional electrical stimulation (FES) has been proposed [8, 9]. The use of FES can induce muscle contractions and utilise the patient’s muscles as actuators. This increases the afferent feedback to the patient, which is associated with increased neuroplasticity [10, 11]. Also, the induced muscle contractions can provide increased calcium concentration in bones and improved cardiovascular benefits [12]. However, with the use of FES the rapid onset of muscle fatigue [13, 14] and the non-linear response of muscles to FES makes the resultant limb motion hard to control [15, 16]. When combined with robotic assistance (Fig. 1), it is possible to overcome these limitations but achieving a seamless triadic collaboration between the user, the robot and FES remains a challenge.
@@
Some of the earlier studies on hybrid systems tested the possibility of combining FES with passive wearable orthoses [17, 18, 19]. However, the passive nature and the dissipative control of these orthoses, implied that the joint torque induced by FES needed to match or exceed the desired joint torque. Thus, even though these passive orthoses proved to be safe to use, their inability to be actively controlled to provide assistive torques that would increase the energy of the system limited their capacity to reduce muscle fatigue.
@@
To address this, more emphasis was given on combining FES with active electromechanical devices. This led to the exploration of a more diverse pool of assistive controllers with the majority of them incorporating adaptive control for either or both the robot and FES. Examples include the studies by Stauffer et al. [20] and Ha et al. [21] where iterative learning control (ILC) was used to adjust the stimulation intensity of FES in order to minimise the interaction torques between the user and the robotic device while performing an ambulatory task. ILC was also used in [22] for both the FES controller and the robotic controller. Del-Ama et al. used a dual state controller that switched between a learning state and a monitoring state to enable the adjustment of the stimulation parameters and the stiffness of the exoskeleton, respectively, in order to respond to muscle fatigue, reduce the interaction between the human and the robot, and maintain an accurate tracking of the desired kinematic trajectory [22].
@@
Studies on hybrid robot-FES cooperation have also been carried out using optimal control and nonlinear control to address the actuation redundancy problem [23, 24, 25]. In [23] the use of non-linear model predictive control (NMPC) was used during the knee extension task to demonstrate good tracking performance as well as disturbance rejection. Similarly, Molazadeh et al. [24] used MPC as a low-level controller to distribute the required assistive torques between a robotic device and FES. These required assistive torques were estimated using a high-level neural network-based iterative learning controller (NNILC) which was designed to learn the unknown system dynamics. This system was shown to be effective in iteratively reducing the trajectory tracking error by reducing the feedback input and increasing the contribution of the learning terms. A slightly different approach was adopted by Alibeji et al. [25] where postural synergies were used to reduce the dimensions of the control variables. Based on the identified postural synergies, the estimated muscle fatigue and a Lyapunov-based stability approach, a non-linear controller was derived which achieved accurate trajectory tracking. However, the controller’s effect on muscle fatigue could not be analysed due to the limited duration of the experiments. Some interesting torque distribution approaches have also been explored in simulation in order to address this actuation redundancy in hybrid controllers but their effectiveness was not validated experimentally [26, 27].
@@ II-A Hybrid Path Control
As our starting point we deploy a Hybrid Path Controller – a controller that follows the principles of Path Control [34], and provides assistance using a wearable robot and FES. This controller involves a reference kinematic path in joint space, ${\bf Q}_{ref}\in\mathbb{R}^{i\times j}$, that describes the desired relationship between the hip and the knee joint angle in the sagittal plane (where $i$ is the number of points in the discretised domain of the reference path and $j$ is the number of degrees of freedom defined by the reference path). Based on this reference path, the reference point, ${\bf q}_{\text{ref}}\in\mathbb{R}^{j}$, is calculated according to the joint angles of the human, ${\bf q}_{\text{act}}\in\mathbb{R}^{j}$ (for this study, the joint angles of the human are approximated as the joint angles of the exoskeleton). Given the joint angles of the human, this reference point is dynamically defined as the point on the reference path where the Euclidean distance between the reference path, ${\bf Q}_{ref}$, and the joint angles of the human, ${\bf q}_{act}$, is at a minimum (Fig. 2). To establish a hierarchy between the human, the FES and the robot, a dead band and an FES band of radius, $r_{db}$ and $r_{fesb}$, respectively, are defined around the reference path. This defines a region where no assistance will be provided to the human (dead band), a region where only assistance from FES will be provided (FES band) and a region where hybrid robot-FES assistance will be provided (region beyond the FES band). The corresponding joint angle error can then be calculated for the FES controller, ${\Delta{\bf q}}_{f}\in\mathbb{R}^{j}$, and the exoskeleton’s controller, ${\Delta{\bf q}}_{e}\in\mathbb{R}^{j}$, as the difference between the reference point and the human’s pose, minus the relevant error tolerance, $r$, formally expressed as:
@@
Based on the tracking error, assistance will be provided by the FES and the robot, in a direction orthogonal to the reference path in order to promote patient-driven therapy without temporal constraints.
@@ II-B Adaptive FES Controller
${\bf u}_{f}={\boldsymbol{\mu}}{\boldsymbol{\gamma}}{\bf K}_{f}\Delta{\bf\tilde{q}}_{f},$ (4)
@@
where ${\bf{u}}_{f}\in\mathbb{R}^{k}$ is the stimulation intensity, $\boldsymbol{\mu}\in\mathbb{R}^{k}$ is a measure of muscle fitness (the inverse of muscle fatigue) ranging from [0,1], $\boldsymbol{\gamma}\in\mathbb{R}^{k}$ is the ILC gain, ${\bf{K}}_{f}\in\mathbb{R}^{k}$ is a constant stiffness for the FES controller, and ${\Delta\tilde{\bf q}}_{f}\in\mathbb{R}^{k}$ is the kinematic error adjusted by its sign and the stimulated muscles to ensure ${\bf u}_{f}\geq 0$, with $k$ being the number of stimulated muscles. ${\bf u}_{thr}$ and ${\bf u}_{sat}$ are the threshold and saturation pulse width for the given stimulation amplitude and frequency, and $r_{\text{fesb}}^{0}$ is the initial radius of the FES band (set to 6 degrees).
The first adaptive component involves the adjustment of the stimulation intensity according to muscle fatigue. That is, muscles whose fatigue increases receive less stimulation. This adaptation occurs at the controller’s frequency while muscle fatigue is estimated using the fitness model [35]:
@@
where ${\bf k}_{1}$ and ${\bf k}_{2}$ are constants that characterise the response of muscles to electrical stimulation as described in [36] including a rise time constant, $T_{rise}$, a fall time constant, $T_{fall}$, and the excitation time constant, $T_{e}$, where $k_{1}=T_{e}*T_{rise/fall}$, and $k_{2}=T_{e}+T_{rise/fall}$.
@@
The second adaptive component of this controller is the stimulation adjustment that occurs iteratively, at every gait cycle, based on the root mean squared error of the user, ${\Delta\bf q}_{f}$. This adaptive component reduces the value of parameter $\gamma$ as the performance of the patient improves until the patient can perform the task without any assistance. For this, a finite state machine (FSM) is also used to differentiate between the swing phase and stance phase where different assistive forces are likely to be appropriate. This is defined as:
${\boldsymbol{\gamma}}^{z+1}_{st}={\boldsymbol{\phi}}_{f}{\boldsymbol{\gamma}}^{z}_{st}+{\boldsymbol{\lambda}}_{f}{\text{rms}({\Delta\bf q}_{f})}^{z}_{st},$ (11)
${\boldsymbol{\gamma}}^{z+1}_{sw}={\boldsymbol{\phi}}_{f}{\boldsymbol{\gamma}}^{z}_{sw}+{\boldsymbol{\lambda}}_{f}{\text{rms}({\Delta\bf q}_{f})}^{z}_{sw},$ (12)
$\lambda_{f}=(1-\phi),$ (13)
where ${\boldsymbol{\phi}_{f}}$ and $\boldsymbol{\lambda}_{f}$ are the forgetting and learning factors of the FES ILC (${\boldsymbol{\phi}}_{f}$ set to 0.95), and $z$ is the number of gait cycles performed.
@@ II-C Adaptive Exoskeleton Controller
where ${\bf K}_{e}$ and ${\bf B}_{e}$ are the joint stiffness and damping of the exoskeleton’s joints, respectively, and ${\bf c}_{cr}$ is the matrix of the critical damping coefficients.
Like the FES controller, the exoskeleton controller also uses a FSM in order to adjust the stiffness of the exoskeleton based on the phase in the gait cycle. On every gait cycle, the root mean squared error of the user, ${\Delta\bf q}_{e}$, is recorded and is used to update the exoskeleton’s stiffness as follows:
${\bf K}^{z+1}_{st}={\boldsymbol{\phi}}_{e}{\bf K}^{z}_{st}+{\boldsymbol{\lambda}}_{e}{\bf K}_{st}^{0}{\text{rms}({\Delta\bf q}_{e})}^{z}_{st},$ (16)
${\bf K}^{z+1}_{sw}={\boldsymbol{\phi}}_{e}{\bf K}^{z}_{sw}+{\boldsymbol{\lambda}}_{e}{\bf K}_{sw}^{0}{\text{rms}({\Delta\bf q}_{e})}^{z}_{sw},$ (17)
$\lambda_{e}=(1-\phi),$ (18)
@@ II-D Adaptive FES Band
The adaptation of the width of the FES band is designed in order to achieve the collaboration between the robot and the FES. By reducing the width of the FES band according to the fitness level of the muscles, robotic assistance will be combined with FES at an earlier stage. This adaptation of the FES band is implemented as:
@@ III-A Subject
The effectiveness of the HAPC was tested both in simulation and experimentally on one healthy volunteer (male, age = 30, weight = 80 kg). The experiment pipeline was approved by the University of Edinburgh, School of Informatics Ethics Committee and the participant provided written consent.
Figure 3: Simulated response of the hybrid controller. (a) Shows the robotic stiffness and assistance provided, (b) the trajectory error and adaptation of parameter $\gamma$, and (c) the muscle fatigue and stimulation provided with the adaptive controller during poor model performance and more accurate performance (change in model behaviour indicated by a grey dotted line).
Figure 4: Comparison of the normalised trajectory error, robotic assistance, FES, and muscle fatigue between the exoskeleton path controller (EPC), the FES path controller (FPC), the hybrid path controller (HPC) and the hybrid adaptive path controller (HAPC).
@@ III-C Fatigue Model Identification
To ensure that the estimated fatigue of muscles reflects the fitness level of the participant, a model calibration was carried out as described in [37, 38]. The feet of the participant were placed in contact with a heavy (immovable) box and the generated muscle force (contact force) was measured using the treadmill’s 6-axis force-torque sensors. A symmetric biphasic signal with constant stimulation frequency of 25 Hz and constant current amplitude of 25 mA was used under isometric conditions (Fig. 1a). Firstly, to identify the minimum stimulation pulse width that results in a visible muscle contraction, $u_{thr}$, and the pulse width above which the generated force does not increase, $u_{sat}$, a series of signals were delivered where the pulse width was increased by 50 $\mu s$ until a pulse width of 800 $\mu s$ was reached. The signal duration was set to 4 s, with a resting period of 20 s. A fatigue test was then carried out using a continuous signal at the saturation intensity for 180 s, followed by a 2-minute recovery test which included pulses of 1 s, with a between-pulse rest of 10 s. The identification of the model parameters was carried out using MATLAB’s optimisation tool fmincon. The identified parameters are presented in Table I.
@@ III-D Simulation Setup
The initial development and testing of the controller was carried out using musculoskeletal modelling. Using the biomechanical modelling software OpenSim [39], a generic lower-limb model, gait1018, was scaled to the dimensions of the participant and was combined with the H3 exoskeleton as described in [33]. An estimate of the voluntary joint torques needed for the tracking of the trajectory was obtained using motion capture as described in [33] and forward dynamics simulations were carried out to predict the tracking error, muscle fatigue and assistance provided to the model. The exoskeleton assistance was simulated using ideal joint actuators while an estimate of the FES-induced torques was obtained based on Hill-type muscle models [40] and the activation dynamics described in section II-B for the monoarticular muscles of both the hip joint and the knee joint.
To observe the controller’s response to a changing human behaviour, simulations were carried out with two different inputs for the human controls. In the first instance, and for the first half of the simulation, a behaviour that corresponds to a high-error trajectory was used, while for the second half, a behaviour that corresponds to a low-error trajectory was used. These trajectories were obtained from our previous study in [33]. Using these model controls, a comparison was also carried out between the hybrid adaptive path controller (HAPC), its non adaptive equivalent (HPC), and the exoskeleton-only (EPC) and FES-only controllers (FPC). In all cases, 64 cycles were simulated, for which the RMS error, RMS of the assistance and mean fatigue were used for the comparison.

### 08 Mar 2024 v2

@@ I INTRODUCTION
To further enhance the outcomes of physiotherapy, the combination of robotic assistance with functional electrical stimulation (FES) has been proposed [8, 9]. The use of FES can induce muscle contractions and utilise the patient’s muscles as actuators. This can provide increased calcium concentration in bones and improved cardiovascular benefits [10]. Also, the afferent feedback to the patient is increased, which is associated with increased neuroplasticity [11, 12]. However, with the use of FES the rapid onset of muscle fatigue [13, 14] and the non-linear response of muscles make the induced motion hard to control [15, 16]. When combined with robotic assistance (Fig. 1), these limitations can be overcome but achieving a seamless triadic collaboration between the user, the robot and FES remains a challenge.
@@
Some of the earlier studies on hybrid systems tested the possibility of combining FES with passive wearable orthoses [17, 18, 19]. However, the passive nature and the dissipative control of these orthoses, implied that the joint torque induced by FES needed to match or exceed the desired joint torque. Thus, even though these passive orthoses proved to be safe, their inability to be actively controlled to provide assistive torques that would increase the energy of the system limited their capacity to reduce muscle fatigue.
@@
To address this, more emphasis was given on combining FES with active electromechanical devices. This led to the exploration of a more diverse pool of assistive controllers with the majority of them incorporating adaptive control for either or both the robot and FES. Examples include the studies by Stauffer et al. [20] and Ha et al. [21] where iterative learning control (ILC) was used to adjust the stimulation intensity of FES in order to minimise the interaction torques between the user and the robot while performing an ambulatory task. ILC was also used in [22] for both the FES controller and the robotic controller. Del-Ama et al. used a dual state controller that switched between a learning state and a monitoring state to enable the adjustment of the stimulation parameters and the stiffness of the exoskeleton, respectively, in order to respond to muscle fatigue, reduce the interaction between the human and the robot, and maintain an accurate tracking of the desired kinematic trajectory [22].
@@
Studies on hybrid robot-FES cooperation have also been carried out using optimal control and nonlinear control to address the actuation redundancy problem [23, 24, 25]. In [23] the use of non-linear model predictive control (NMPC) was used during a knee extension task and demonstrated good tracking performance and disturbance rejection. Similarly, in [24] MPC was used to distribute the assistive forces between a robotic device and FES. These assistive forces were estimated using a neural network-based ILC (NNILC) which was designed to learn the unknown system dynamics. This system was shown to be effective in reducing the tracking error by reducing the feedback input and increasing the contribution of the learning terms. A slightly different approach was adopted by Alibeji et al. [25] where postural synergies were used to reduce the number of control variables. Based on the identified postural synergies, the estimated muscle fatigue and a Lyapunov-based stability approach, a non-linear controller was derived which achieved accurate trajectory tracking. However, the controller’s effect on muscle fatigue could not be analysed due to the limited duration of the experiments. Some interesting torque distribution methods have also been explored in simulation in order to address this actuation redundancy in hybrid control but their effectiveness was not validated experimentally [26, 27].
@@ II-A Hybrid Path Control
As our starting point we deploy a Hybrid Path Controller – a controller that follows the principles of Path Control [34], and provides assistance using a wearable robot and FES. This controller involves a reference kinematic path in joint space, ${\bf Q}_{ref}\in\mathbb{R}^{i\times j}$, that describes the desired relationship between the hip and the knee joint angle in the sagittal plane (where $i$ is the number of points in the discretised domain of the reference path and $j$ is the number of degrees of freedom defined by the reference path). Based on this reference path, the reference point, ${\bf q}_{\text{ref}}\in\mathbb{R}^{j}$, is calculated according to the joint angles of the human, ${\bf q}_{\text{act}}\in\mathbb{R}^{j}$ (for this study, the joint angles of the human are approximated as the joint angles of the exoskeleton). Given the joint angles of the human, this reference point is dynamically defined as the point on the reference path where the Euclidean distance between the reference path, ${\bf Q}_{ref}$, and the joint angles of the human, ${\bf q}_{act}$, is at a minimum (Fig. 2). To establish a hierarchy between the human, the FES and the robot, a dead band and a FES band of radius, $r_{db}$ and $r_{fesb}$, respectively, are defined around the reference path. This defines a region where no assistance is provided to the human (dead band), a region where only assistance from FES is provided (FES band) and a region where hybrid robot-FES assistance is provided (hybrid band). The corresponding joint angle error can then be calculated for the FES controller, ${\Delta{\bf q}}_{f}\in\mathbb{R}^{j}$, and the exoskeleton’s controller, ${\Delta{\bf q}}_{e}\in\mathbb{R}^{j}$, as the difference between the reference point and the human’s pose, minus the relevant error tolerance, $r$, formally expressed as:
@@
Based on the tracking error, assistance is provided by the FES and the robot, in a direction orthogonal to the reference path in order to promote patient-driven therapy without temporal constraints. For this study the reference path was defined based on the recorded motion of a healthy subject and was adjusted to a path with a less pronounced loading response.
@@ II-B Adaptive FES Controller
${\bf u}_{f}={\boldsymbol{\mu}}{\boldsymbol{\gamma}}{\bf K}_{f}\Delta{\bf\overline{q}}_{f},$ (4)
@@
where ${\bf{u}}_{f}\in\mathbb{R}^{k}$ is the stimulation intensity, $\boldsymbol{\mu}\in\mathbb{R}^{k}$ is a measure of muscle fitness (the inverse of muscle fatigue) ranging from [0,1], $\boldsymbol{\gamma}\in\mathbb{R}^{k}$ is the ILC gain ranging from [0,1], ${\bf{K}}_{f}\in\mathbb{R}^{k}$ is a constant stiffness, and ${\Delta\overline{\bf q}}_{f}\in\mathbb{R}^{k}$ is the kinematic error adjusted by its sign and the stimulated muscles to ensure ${\bf u}_{f}\geq 0$, with $k$ being the number of stimulated muscles. ${\bf u}_{thr}$ and ${\bf u}_{sat}$ are the threshold and saturation pulse width for the given stimulation amplitude and frequency, and $r_{\text{fesb}}^{0}$ is the initial radius of the FES band (set to 6°).
The first adaptive component involves the adjustment of the stimulation intensity according to muscle fatigue. That is, muscles whose fatigue increases receive less stimulation. To estimate muscle fatigue the following model is used [35]:
@@
where ${\bf k}_{1}$ and ${\bf k}_{2}$ are constants that characterise the response of muscles to electrical stimulation as described in [36] including a rise time constant, $T_{rise}$, a fall time constant, $T_{fall}$, and the excitation time constant, $T_{e}$, where $k_{1}=T_{e}*T$, and $k_{2}=T_{e}+T$ for $T=T_{rise}$, if $e>a$ and $T=T_{fall}$, otherwise.
@@
The second adaptive component of this controller is the stimulation adjustment that occurs iteratively, at every gait cycle, based on the root mean squared error of the user. This adaptive component reduces the value of parameter $\gamma$ as the performance of the patient improves until the patient can perform the task without any assistance. For this, a finite state machine (FSM) is used to differentiate between the swing phase (SW) and stance phase (ST) where different assistive forces may be appropriate. This is defined as:
${\boldsymbol{\gamma}}^{z+1}_{st}={\boldsymbol{\phi}}_{f}{\boldsymbol{\gamma}}^{z}_{st}+{\boldsymbol{\lambda}}_{f}{\text{rms}({\Delta\bf\tilde{q}}_{f})}^{z}_{st},$ (11)
${\boldsymbol{\gamma}}^{z+1}_{sw}={\boldsymbol{\phi}}_{f}{\boldsymbol{\gamma}}^{z}_{sw}+{\boldsymbol{\lambda}}_{f}{\text{rms}({\Delta\bf\tilde{q}}_{f})}^{z}_{sw},$ (12)
$\lambda_{f}=(1-\phi_{f}),$ (13)
where ${\boldsymbol{\phi}_{f}}$ and $\boldsymbol{\lambda}_{f}$ are the forgetting and learning factors of the FES ILC (${\boldsymbol{\phi}}_{f}$ set to 0.95), ${\Delta\bf\tilde{q}}_{f}$ is the error normalised by $r_{db}$ (set to 2°) and $z$ is the number of gait cycles performed.
@@ II-C Adaptive Exoskeleton Controller
where ${\bf K}_{e}$ and ${\bf B}_{e}$ are the stiffness and damping of the exoskeleton’s joints, respectively, and ${\bf c}_{cr}$ is the matrix of the critical damping coefficients.
Like the FES controller, the exoskeleton controller also uses a FSM in order to adjust the stiffness of the exoskeleton based on the phase in the gait cycle. On every gait cycle, the root mean squared error of the user is recorded and is used to update the exoskeleton’s stiffness as follows:
${\bf K}^{z+1}_{st}={\boldsymbol{\phi}}_{e}{\bf K}^{z}_{st}+{\boldsymbol{\lambda}}_{e}{\bf K}_{e}^{0}{\text{rms}({\Delta\bf\tilde{q}}_{e})}^{z}_{st},$ (16)
${\bf K}^{z+1}_{sw}={\boldsymbol{\phi}}_{e}{\bf K}^{z}_{sw}+{\boldsymbol{\lambda}}_{e}{\bf K}_{e}^{0}{\text{rms}({\Delta\bf\tilde{q}}_{e})}^{z}_{sw},$ (17)
$\lambda_{e}=(1-\phi_{e}),$ (18)
@@ II-D Adaptive FES Band
The adaptation of the width of the FES band is designed in order to achieve the collaboration between the robot and the FES. By reducing the width of the FES band according to the fitness level of the muscles, robotic assistance is combined with FES at an earlier stage. This adaptation of the FES band is implemented as:
@@ III-A Subject
The effectiveness of the HAPC was tested both in simulation and experimentally on one healthy volunteer (male, age = 30, weight = 80 kg). The experiment pipeline was approved by the University of Edinburgh, School of Informatics Ethics Committee (ID 2021/46920) and the participant provided written consent.
Figure 3: Simulated response of the hybrid controller. (a) Shows the robotic stiffness and assistance provided, (b) the trajectory error and adaptation of parameter $\gamma$, and (c) the muscle fatigue and stimulation provided with the adaptive controller during poor model performance (before grey dotted line) and more accurate performance (after grey dotted line).
Figure 4: Comparison of the normalised trajectory error, robotic assistance, FES, and muscle fatigue between the exoskeleton controller (EPC), the FES controller (FPC), the hybrid controller (HPC) and the hybrid adaptive controller (HAPC).
@@ III-C Fatigue Model Identification
To ensure that the estimated fatigue of muscles reflects the fitness level of the participant, a model calibration was carried out as described in [37, 38]. The feet of the participant were placed in contact with a heavy (immovable) box and the generated muscle force (contact force) was measured using the treadmill’s force sensors. A symmetric biphasic signal with constant stimulation frequency of 25 Hz and current amplitude of 25 mA was used under isometric conditions (Fig. 1a). Firstly, to identify the minimum stimulation pulse width that results in a visible muscle contraction, $u_{thr}$, and the pulse width above which the generated force saturates, $u_{sat}$, a series of signals were delivered at 50 $\mu s$ increments until a pulse width of 800 $\mu s$ was reached. The signal duration was set to 4 s, with a resting period of 20 s. A fatigue test was then carried out using a continuous signal at the saturation intensity for 180 s, followed by a 2-minute recovery test which included pulses of 1 s, with a between-pulse rest of 10 s. The identification of the model parameters was carried out using MATLAB’s optimisation tool fmincon. The identified parameters are presented in Table I.
@@ III-D Simulation Setup
The initial development and testing of the controller was carried out using musculoskeletal modelling. Using the biomechanical modelling software OpenSim [39], a generic lower-limb model, gait1018, was scaled to the dimensions of the participant and was combined with the H3 exoskeleton as described in [33]. An estimate of the voluntary joint torques needed for the tracking of the trajectory was obtained using motion capture and forward dynamics simulations were carried out to predict the tracking error, muscle fatigue and assistance provided to the model [33]. The exoskeleton assistance was simulated using ideal joint actuators while an estimate of the FES-induced torques was obtained based on Hill-type muscle models [40] and the activation dynamics described in section II-B for the monoarticular muscles of both the hip joint and the knee joint.
To observe the controller’s response to a changing human behaviour, simulations were carried out with two different inputs for the human controls. For the first half of the simulation, a behaviour that corresponds to a high-error trajectory was used, while for the second half, a behaviour that corresponds to a low-error trajectory was used. These trajectories were obtained from our previous study in [33]. Using these model controls, a comparison was also carried out between the hybrid adaptive path controller (HAPC), its non adaptive equivalent (HPC), and the exoskeleton-only (EPC) and FES-only controllers (FPC). In all cases, 64 cycles were simulated, for which the RMS error, RMS of the assistance and mean fatigue were used for comparison.